OpenStudy (anonymous):

Given U = {22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32}, A = {22, 24, 26, 28, 30}, and B = {23, 24, 25, 28, 29}. Find A′ ∩ B′. Do not skip to the answer. Show your work, finding A’, then B’ then find their intersection

OpenStudy (anonymous):

U is the Universe or all possible values. A is a subset of U, and A' is also a subset and is all elements of U that are not in A. Similar for B'. Once you have A' and B', the intersection is just thos elements that are in both A' and B'. So, start by listing A' and B' and you will have no problem.

OpenStudy (anonymous):

Im so confused :(

OpenStudy (anonymous):

I can't give you any more without just giving you the answer.

OpenStudy (anonymous):

And I don't want that because thats not how I am going to learn... I appreciate your help and I will look further into it and try to figure it out...You did give me a start.. thanks :)

OpenStudy (anonymous):

Don't give up so easily! Tell me where you are confused and maybe I can guide you into getting the answer step-by-step.

OpenStudy (anonymous):

U, A, A', B, and B' are all sets. A + A' = U. So does B + B' = U. A' contains everything that is in U that is NOT also in A. That's a good start: find out what is in A'.

OpenStudy (anonymous):

Oh ok.. I think I have this... does this look correct? One sec let me type it out

OpenStudy (anonymous):

Hint: there are 11 elements in U. There are 5 elements in A. There will be 6 elements in A' because 11 - 5 = 6. That's a handy thing to do when checking if you are doing your set work correctly.

OpenStudy (anonymous):

Ok so this is how I would post my answer to "show my work" and knowledge of the steps.... Given that the universe (U) is {22…32}, and A={22, 24, 26, 28, 30}, and B={23…25, 28, 29}, find A'∩B': First, to find A', take all elements of U that don't have corresponding elements in A: {23, 25, 27, 29, 31, 32}. Next, to find B', take all elements of U that don't have corresponding elements in B: {22, 26, 27, 30…32}. Last, find the intersection of those two intermediate product sets: 23 is in A'. In B'? No. 25 is in A' only, so no. 27 is in both, so yes. 29 is in A' only, so no. 31 and 32 are both in A' and in B', so yes. The final set is {27, 31, 32}.

OpenStudy (anonymous):

You've got it and you did it very well. By your answer, which is well-stated, you must have gotten rid of any confusion, because you certainly understand it well now! Good job!

OpenStudy (anonymous):

Thank you so much!! I have good knowledge of this lesson, I just think I was confused at what the question was asking me for some reason.... Thanks so much for your clarification!! Have a good one!!!

OpenStudy (anonymous):

good luck in all your future math and everything else!

OpenStudy (anonymous):

Thanks!!